9.Rapid-Lube provides oil changes and lubes. The estimated number of oil changes for April and May are 3,600 and 4,000. Each oil change takes 12 minutes of direct labour. The wage rate is $10 per hour. Overhead is $3,700 per month and $2 per oil change.

A.Budgeted direct labour for April is $__________________.

B.Budgeted direct labour for May is $__________________.

C.Budgeted overhead for April is $__________________.

D.Budgeted overhead for May is $__________________.

10.Borland Company makes backpacks. Its production budget for two months is:

Month Budgeted production in units

June35,000                       

July50,000

Borland uses two types of labour to make the backpacks: cutting labour and sewing labour. Each backpack requires 6 minutes, on average, of cutting labour. Each backpack requires 24 minutes of sewing labour.

Borland has fixed overhead of $4,400 per month and variable overhead of $3 per direct labour hour.

A.How many hours of cutting labour are budgeted for July?

B.How many hours of sewing labour are budgeted for July?

C.What is the total amount of budgeted direct labour hours for July?

D.What is the budgeted total overhead for the month of July?
Answer #1

9) A Total oil changes in April = 3600

So, total time of labor that will be required if each oil change takes 12 minutes of direct labor =

3600*12 = 43200 minutes.

Now the budget for direct labor in April will be 43200*10 = $432000.

B) Total oil changes in April = 4000

So, total time of labor that will be required if each oil change takes 12 minutes of direct labor =

4000*12 = 48000 minutes.

Now the budget for direct labor in April will be 48000*10 = $480000.

C) Budgeted overhead for April = 3700 + 2*3600 = $10900.

D) Budgeted overhead for May = 3700 + 2*4000 = $11700.